OpenStudy (anonymous):

State whether the triangles are similar. If so, write a similarity statement and the postulate or theorem you used. pic and medal :)

OpenStudy (anonymous):

what are the triangles?

OpenStudy (anonymous):



OpenStudy (anonymous):

i just did this today acutally

OpenStudy (anonymous):

can u plz do it again? its my last thing for this assignment for math..

OpenStudy (anonymous):

yep! im on it!

OpenStudy (anonymous):

or at least help me to do it myself..thanks, soooo much :)

OpenStudy (anonymous):

30/10 = 3/1 3=3 so, corresponding sides are in proportion

OpenStudy (anonymous):

damn, i knew that..i just didnt know how to prove it, cuz we need to explain it..wow, i feel stupid, lol..thanks :P

OpenStudy (anonymous):

kk

OpenStudy (anonymous):